OpenStudy (anonymous):

find all the zeros of the equations. please show work x^3-5x^2+6x=0 x^4-8x^2-9=0 x^3-125=0

jimthompson5910 (jim_thompson5910):

x^3-5x^2+6x=0 x(x^2 - 5x + 6) = 0 x(x - 3)(x - 2) = 0 I'll let you finish up

jimthompson5910 (jim_thompson5910):

x^4-8x^2-9=0 z^2-8z-9=0 ... let z = x^2, so z^2 = x^4 (z - 9)(z - 1) = 0 (x^2 - 9)(x^2 - 1) = 0 (x-3)(x+3)(x-1)(x+1) = 0 again I'll stop here to let you finish

jimthompson5910 (jim_thompson5910):

x^3-125=0 x^3-5^3=0 (x-5)(x^2 + 5x + 25) = 0 ... use the difference of cubes factoring rule stopping here

OpenStudy (anonymous):

what would i have to do to finsh it up?

jimthompson5910 (jim_thompson5910):

you would use the zero product property to go from x(x - 3)(x - 2) = 0 to x = 0, x-3=0 or x-2=0 then you would solve each equation for x

OpenStudy (anonymous):

im not sure with the zero product proterty

jimthompson5910 (jim_thompson5910):

the zero product property is the idea that if A*B = 0 then either A = 0 or B = 0 (or both could be zero)

OpenStudy (anonymous):

what would i need to do to solve for x?

jimthompson5910 (jim_thompson5910):

x(x - 3)(x - 2) = 0 x = 0, x-3=0 or x-2=0 x=0, x=3, or x = 2

jimthompson5910 (jim_thompson5910):

see how I'm getting this?

OpenStudy (anonymous):

I really appreciate your help. I am really trying hard to understand, but it is so confusing.

jimthompson5910 (jim_thompson5910):

basically if I told you two numbers multiply to 0 then at least one of them must be zero

jimthompson5910 (jim_thompson5910):

so that's how something like x times (x-3) = 0 turns into x = 0 or x-3 = 0

OpenStudy (anonymous):

ok thank you for your help

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

could you help me with another problem please?

OpenStudy (anonymous):

Determine the number and type of complex solutions and possible real solutions for 2x^4+x^2-x+6=0?

jimthompson5910 (jim_thompson5910):

have you ever heard of descartes rule of signs?

OpenStudy (anonymous):

no i have not, sorry

jimthompson5910 (jim_thompson5910):

ok have a look at this page http://www.purplemath.com/modules/drofsign.htm and tell me if it helps or not

OpenStudy (anonymous):

i read the page and i really didn't understand anything at all

jimthompson5910 (jim_thompson5910):

ok I'll try my best to explain the idea

jimthompson5910 (jim_thompson5910):

2x^4+x^2-x+6=0 is the same as +2x^4+1x^2-1x+6=0 notice from the first term +2x^4 to the second term +1x^2, the coefficients are +2 and +1 respectively

OpenStudy (anonymous):

thank you so much, im failing my alg class

jimthompson5910 (jim_thompson5910):

so there is NO sign change in the coefficients

jimthompson5910 (jim_thompson5910):

However... the jump from 1x^2 to -1x has a sign change happening since you go from a positive coefficient (+1) to a negative coefficient (-1)

jimthompson5910 (jim_thompson5910):

see the sign change I'm pointing out?

OpenStudy (anonymous):

yeah i see the sign change

jimthompson5910 (jim_thompson5910):

ok great there is another sign change: as you go from -1x to +6, there is a sign change from negative to positive

jimthompson5910 (jim_thompson5910):

in total, there are 2 sign changes

jimthompson5910 (jim_thompson5910):

which means that there are a maximum of two positive roots

jimthompson5910 (jim_thompson5910):

I'm just using the rule described on that page I sent you Rule: if there are n sign changes, then there are a maximum of n positive roots

jimthompson5910 (jim_thompson5910):

with me so far?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

ok let f(x) = 2x^4+x^2-x+6 we need to find f(-x) now f(x) = 2x^4+x^2-x+6 f(-x) = 2(-x)^4+(-x)^2-(-x)+6 ... replace every 'x' with '-x' f(-x) = 2x^4+x^2+x+6 ... simplify Notice how there are NO sign changes in f(-x) = 2x^4+x^2+x+6 since everything is positive and there are no negative terms

jimthompson5910 (jim_thompson5910):

so this means that there are NO negative real roots

jimthompson5910 (jim_thompson5910):

to sum up so far f(x) = 2x^4+x^2-x+6 has at max two real positive roots f(x) = 2x^4+x^2-x+6 has NO real negative roots

jimthompson5910 (jim_thompson5910):

Now if f(x) = 2x^4+x^2-x+6 had 2 real positive roots, then you would need 2 more complex roots (because you need 4 roots total) if f(x) = 2x^4+x^2-x+6 had 1 real positive root, then you would need 3 more complex roots (because you need 4 roots total)...but complex roots ALWAYS come in pairs...so this scenario isn't possible if f(x) = 2x^4+x^2-x+6 had 0 real positive roots, then you would need 4 more complex roots (because you need 4 roots total)

jimthompson5910 (jim_thompson5910):

it's definitely a lot to take in, so read it over and ask about anything that stumps you

jimthompson5910 (jim_thompson5910):

I guess you could sum it up like this f(x) = 2x^4+x^2-x+6 has at max two real positive roots f(x) = 2x^4+x^2-x+6 has NO real negative roots f(x) = 2x^4+x^2-x+6 has at most 4 complex roots The number of real and complex roots must add to 4

OpenStudy (anonymous):

thank you so much

jimthompson5910 (jim_thompson5910):

sure thing